# Non-Gravitational Forces in Planetary Systems

### v1

@@ 7 Lorentz Force
A charged particle interacts with a moving magnetic field through the Lorentz force, $\mathcal{F}_{L}$. For simplicity, we ignore the fact that $\mathcal{F}_{L}$ is a vector acting perpendicular to both the velocity of the particle and the directtion of the field, and we consider only the magnitude of the force, $\mathcal{F}_{L}=BqV$. Here $B$ is the magnetic flux density, $q$ is the charge on the particle and $V$ is the velocity of the particle with respect to the field.
@@
for the charge on a particle of radius $a$ when its potential, $U$, is known. Numerous effects contribute to the charging of dust particles in the solar system, but the dominant effect is photoionization from solar ultraviolet, which leads to $U\sim$ 5 to 10 Volts (Wyatt, 1969). This potential can vary depending on fluctuations in the solar ultraviolet flux, itself a function of the solar magnetic cycle. The potential, $U$, depends mainly on the photon energy and the ionization threshold of the material, and is therefore approximately independent of distance from the Sun. However, the charging time increases with $r_{H}^{2}$, and so is $\sim 10^{3}$ times longer in the Kuiper belt at 30 au than near the Earth at 1 au, elevating the relative importance of leakage currents that ultimately can limit the accumulated charge. Here we take $U$ = 10 Volts independent of distance.
@@ 7.1 Examples
    Three exceptions are the long-period comets C/1995 O1 (Hale-Bopp), C/2014 B1 (Schwartz) and C/2010 U3 (Boattini). C/1995 O1(Hale-Bopp) was observed at $r_{au}\sim$ 20 and reported to show a coma of small particles with $a_{\mu}\sim$ 1 (Kramer et al., 2014). The coma could not be well matched by models using only radiation pressure. With the above values, Equation 32 gives $\beta_{L}/\beta\sim$ 4, consistent with a Lorentz force dominated morphology. C/2014 B1 (Schwartz) was dominated by quite large ($a_{\mu}>$ 100) particles at $r_{au}\sim$ 10 (Jewitt et al., 2019). It has $\beta_{L}/\beta\lesssim 0.02$ by Equation 32, and showed no effect from Lorentz force. Observations of C/2010 U3 (Boattini) at larger distance ($r_{au}\sim$ 20 au) showed smaller particles ($a_{\mu}$ = 10) (Hui et al., 2019). Equation 32 gives $\beta_{L}/\beta\sim$ 0.4, consistent with a larger but still not dominant influence of the Lorentz force. In the Kuiper belt, at $r_{H}$ = 40 au, the motion of all particles smaller than about 8 $\mu$m, including the micron-sized particles detected by the New Horizons particle counter (Doner et al., 2024), should be affected by the Lorentz force.
@@
  • Dust in Planetary Magnetospheres: The giant planets sustain strong dynamo-generated magnetic fields and provide numerous examples where magnetic forces are important. For example, the equatorial surface field of Jupiter is $B_{J}\sim$ 400 $\mu$T, about 105 times stronger than the solar wind field near Earth, and magnetic forces are of correspondingly greater significance. Strong magnetospheric fields are able to retain plasma that would otherwise be quickly stripped away by the solar wind, creating a local gas environment that is distinct from the open interplanetary medium. The magnetospheres of Jupiter and Saturn are filled with dense plasma ($\sim 10^{8}$ to 109 m-3, or 102 to 103 times the solar wind density) from the outgassing satellites Io and Enceladus, respectively. Dust near these satellites is charged negatively by contact with the plasma (because the electrons travel faster in thermal equilibrium and so deliver a larger charging current than the heavier, slower ions). The charge reverses to positive values at larger distances where the effects of plasma charging are overcome by those of UV photoelectron loss (Spahn et al., 2019). Charged dust particles are accelerated by the planetary field, which sweeps past at the corotation speed. One consequence of magnetospheric dust interactions is that Jupiter (Grun et al., 1993) and Saturn (Kempf et al., 2005) eject streams of nanometer sized particles (mass $\sim 10^{-24}$ kg, magnetic $\beta_{L}\gg$ 1) at speeds above the gravitational escape speed from either planet (Hsu et al., 2012). Another is that the B ring of Saturn displays transient, 104 km scale, quasi-radial dust structures known as spokes, whose abundance is seasonally modulated by the ring illumination but also by Saturn’s rotating magnetic field (Smith et al. (1981), Porco & Danielson (1982)). The origin of spokes is still not fully understood, but it is clear that they consist of micron-sized dust particles briefly launched above the Saturn ring plane. The motion of spoke dust reflects the combined influence of local electrostatic and magnetic forces, as well as planetary gravity (Horányi et al., 2004).
@@ 8 Torques
Most asteroids and cometary nuclei are irregular in shape. On such objects, the non-gravitational forces described above generally do not pass through the center of mass, resulting in a torque. Left unchecked, the torque will drive inexorably towards rotational instability, where the centripetal forces exceed the gravitational and cohesive forces holding the body together. For a strengthless sphere of density $\rho$, this occurs at the critical rotation period
$\tau_{crit}=\left(\frac{3\pi}{G\rho}\right)^{1/2}.$ (33)
For example, substituting $\rho=10^{3}$ kg m-3 gives $\tau_{crit}$ = 3.3 hour. A strengthless ellipsoid having equatorial axes $a$ = $b$ and polar axis $c\leq a$ has critical period
$\tau_{crit}=\left(\frac{3\pi}{G\rho}\right)^{1/2}\left(\frac{a}{c}\right)^{1/2}$ (34)
so that a highly oblate body of the same density but with $a/c$ = 2 would have a critical period 21/2 times longer, or 4.7 hours. Aspherical and less dense bodies can have critical periods longer than given by Equation 33, while higher densities and material cohesion lead to shorter critical periods. A majority of the measured rotation periods of solar system bodies are within a factor of a few of $\tau_{crit}$ (Warner et al., 2009), although it must be admitted that observational biases against the measurement of periods either much shorter or much longer than a few hours are strong in most existing datasets.
@@ 8.1 Sublimation Torque
The magnitude of the sublimation torque is equal to the momentum lost per second in sublimated material multiplied by the moment arm, which we write as $k_{T}a$, where $a$ is the radius and 0 $\leq k_{T}\leq 1$ is a dimensionless multiplier. $k_{T}$ = 0 corresponds to perfectly isotropic sublimation with no net torque. $k_{T}$ = 1 corresponds to perfectly collimated ejection tangential to the surface of the body. Then, with a mass loss rate $\mu m_{H}Q_{g}$ [kg s-1] and an outflow speed $V_{th}$, the magnitude of the torque is $T=\mu m_{H}Q_{g}k_{T}V_{th}a$. The timescale for the torque to change the spin is then (Jewitt, 2021)
@@
Measured spin changes in short-period comets allow $\tau_{s}$ to be estimated. With additional observational constraints on $\rho$, $a$, $Q_{g}$ and $P$, and with the use of Equation 35, the dimensionless moment arm $k_{T}$ can be estimated. The median value is $k_{T}$ = 0.007 (meaning that only 0.7% of the outflow momentum is used to torque the nucleus; (Jewitt, 2021)). Even this tiny fraction is sufficient to quickly modify the spins of small comets.
@@
A necessary condition for Equation 35 to remain valid is the persistence of ice at or near the physical surface of the nucleus for timescales $\geq\tau_{s}$. Near surface ice persists because the surface itself moves down as the ice sublimates, such that fresh ice is continually excavated. The particular problem is that refractory particles too large to be ejected by gas drag should eventually clog the surface, producing a ”rubble mantle” and inhibiting or shutting down further sublimation. How this works in detail is not known, even after several years of in-situ investigation of the nucleus of 67P/Churyumov-Gerasimenko by the Rosetta spacecraft (Attree et al., 2023). Orbital evolution may play a role, particularly when the perihelion distance migrates to smaller values, leading to higher temperatures. In the case of the active asteroids, ice is exposed only intermittently, perhaps in response to occasional impacts that clear a surface refractory mantle. The duty cycle (ratio of the time spent in sublimation to the total elapsed time) is less than $10^{-4}$ or even $10^{-5}$, allowing ice to persist for long times but rendering Equation 36 inapplicable to these objects.
@@ 8.2 YORP Torque
A major limitation to the order-of-magnitude estimates of the torque timescales, both for sublimation (Equation 36) and for YORP (Equation 38), concerns the stability of the torque. In comets, the magnitude and direction of the outgassing torque depend on the surface distribution and angular dependence of the outgassing. As the surface erodes, we expect the spatial and angular distribution of outgassing sources to change, altering the torque. For most comets, we possess few or no observational constraints on the areal distribution of sources or their evolution. Indeed, in-situ measurements from 67P/Churyumov-Gerasimenko show the difficulty in modeling this process even given the most detailed data (Attree et al., 2023). Likewise, the YORP torque is highly sensitive to the surface shape and texture and can change in magnitude and even direction in response to minor surface changes (Statler, 2009). Empirical but indirect evidence for this comes from asteroid (162173) Ryugu ($a\sim$ 0.5 km) which has the characteristic diamond shape indicative of rotational instability but a current rotation period near 7.6 hours. The rotation of Ryugu may have slowed in response to a changing YORP torque, leaving its equatorial ridge as evidence of its previously rapid spin. Spin evolution in the presence of a changing torque may be more akin to a random walk process than to the steady change implicit in Equations 36 and 38 (Bottke et al. (2015), Statler (2015)). The relevant spin-change timescales would then be much longer than estimated here. Unfortunately, we possess too little information to address this problem with any confidence.
@@ 8.3 Examples:
    Figure 17: Two active asteroids showing episodic ejections likely due to YORP-driven rotational instability. Each tail corresponds to a particular synchrone; the tail position angle is a measure of the date of ejection. 311P/Panstarrs (top) experienced nine ejections over about nine months (six visible on 2013 Sep 10 plus three later) (Jewitt et al., 2013) while (6478) Gault displayed three tails in 2019 (Luu et al., 2021).
@@ 8.4 Summary
where $\sigma=5.67\times 10^{-8}$ W m-2 K-4 is the Stefan-Boltzmann radiation constant and we assume emissivity $\varepsilon$ = 1. The order-of-magnitude cooling timescale is $\tau_{c}\sim H/(dH/dt)$, so that we can use Equations 3 and 4 to write the thermal parameter (c.f. Spencer et al. (1989))
$\frac{\tau_{c}}{P}=\frac{\rho c_{p}}{\sigma T^{3}}\left(\frac{\kappa}{P}\right)^{1/2}.$ (5)
Asteroids with very rapid rotation ($\tau_{c}/P\gg$ 1) retain their heat through the night and should remain nearly longitudinally isothermal, producing little net diurnal Yarkovsky force. Similarly, asteroids with $\tau_{c}/P\ll$ 1 will experience little diurnal Yarkovsky drift because rotation is too slow to carry the diurnal temperature maximum away from the meridian line. On the other hand, asteroids with $\tau_{c}/P\lesssim$ 1 should retain heat into the afternoon but lose it by the morning and so will experience a net force from asymmetric radiation.
As a rough example, consider an asteroid with a very porous surface regolith having $\kappa$ = 10-9 m2 s-1 and with nominal $\rho=10^{3}$ kg m-3, $c_{p}=10^{3}$ J K-1 kg-1, $P=2\times 10^{4}$ s (i.e., about 6 hours) and orbiting near 1 au, where $T\sim$ 300 K. Substitution gives skin depth $\ell\sim$ 5 mm, $\tau_{c}/P\sim$ 0.15, and the peak heat of noon will be retained for about 1/6th of a rotation, corresponding to a lag angle $\theta\sim$60°. The Yarkovsky force on a small asteroid with these parameters is likely to be significant. By comparison, an asteroid having the same properties but consisting of solid rock, for which $\kappa=10^{-6}$ m2 s-1 would have $\ell\sim$ 15 cm, $\tau_{c}/P\sim$ 4.5 and, by virtue of taking several rotations in order to cool, would experience minimal diurnal temperature variation and reduced Yarkovsky acceleration.
@@
I thank Marco Fenucci for providing a digital table of his Yarkovsky drift measurements, Pedro Lacerda, Jane Luu, Darryl Seligman, David Vokrouhlicky, Emerson Whittaker and an anonymous referee for comments.

### v2

@@ 7 Lorentz Force
A charged particle interacts with a moving magnetic field through the Lorentz force, $\mathcal{F}_{L}$. For simplicity, we ignore the fact that $\mathcal{F}_{L}$ is a vector acting perpendicular to both the velocity of the particle and the direction of the field, and we consider only the magnitude of the force, $\mathcal{F}_{L}=BqV$. Here $B$ is the magnetic flux density, $q$ is the charge on the particle and $V$ is the velocity of the particle with respect to the field.
@@
for the charge on a particle of radius $a$ when its potential, $U$, is known. Numerous effects contribute to the charging of dust particles in the solar system, but the dominant effect is photoionization from solar ultraviolet, which leads to $U\sim$ 5 to 10 Volts (Wyatt, 1969) (corresponding to photon wavelengths $\sim$1000Å to 2000Å). This potential can vary depending on fluctuations in the solar ultraviolet flux, itself a function of the $\sim$22 year solar magnetic cycle. The potential, $U$, depends on the photon energy and the ionization threshold of the material, and is therefore approximately independent of distance from the Sun. However, the photon flux varies as $r_{H}^{-2}$ and so the charging time increases as $r_{H}^{2}$; it is $\sim 10^{3}$ times longer in the Kuiper belt at 30 au than near the Earth at 1 au, elevating the relative importance of leakage currents that ultimately can limit the accumulated charge. In the following we take $U$ = 10 Volts independent of distance.
@@ 7.1 Examples
    Three exceptions are the long-period comets C/2014 B1 (Schwartz), C/2010 U3 (Boattini) and C/1995 O1 (Hale-Bopp). C/2014 B1 (Schwartz) was dominated by quite large ($a_{\mu}>$ 100) particles at $r_{au}\sim$ 10 (Jewitt et al., 2019). It has $\beta_{L}/\beta\lesssim 0.02$ by Equation 32, and showed no effect from Lorentz force. Observations of C/2010 U3 (Boattini) at larger distance ($r_{au}\sim$ 20 au) showed smaller particles ($a_{\mu}$ = 10) (Hui et al., 2019). Equation 32 gives $\beta_{L}/\beta\sim$ 0.4, consistent with a larger but still not dominant influence of the Lorentz force. C/1995 O1 (Hale-Bopp) was observed at $r_{au}\sim$ 20 and reported to show a coma of small particles with $a_{\mu}\sim$ 1 (Kramer et al., 2014). The coma could not be well matched by models using only radiation pressure. With the above values, Equation 32 gives $\beta_{L}/\beta\sim$ 4, consistent with a Lorentz force dominated morphology. In the Kuiper belt, at $r_{H}$ = 40 au, the motion of all particles smaller than about 8 $\mu$m, including the micron-sized particles detected by the New Horizons particle counter (Doner et al., 2024), should be affected by the Lorentz force.
@@
  • Dust in Planetary Magnetospheres: The giant planets sustain strong dynamo-generated magnetic fields and provide numerous examples where magnetic forces are important. For example, the equatorial surface field of Jupiter is $B_{J}\sim$ 400 $\mu$T, about 105 times stronger than the solar wind field near Earth, and magnetic forces are of correspondingly greater significance. Strong magnetospheric fields are able to retain plasma that would otherwise be quickly stripped away by the solar wind, creating a local gas environment that is distinct from the open interplanetary medium. The magnetospheres of Jupiter and Saturn are filled with dense plasma ($\sim 10^{8}$ to 109 m-3, or 102 to 103 times the solar wind density) from the outgassing satellites Io and Enceladus, respectively. Dust near these satellites is charged negatively by contact with the plasma (because the electrons travel faster in thermal equilibrium and so deliver a larger charging current than the heavier, slower ions). The charge reverses to positive values at larger distances where the effects of plasma charging are overcome by those of UV photoelectron loss (Spahn et al., 2019). Charged dust particles are accelerated by the planetary field, which sweeps past at the corotation speed. One consequence of magnetospheric dust interactions is that Jupiter (Grun et al., 1993) and Saturn (Kempf et al., 2005) eject streams of nanometer sized particles (mass $\sim 10^{-24}$ kg, magnetic $\beta_{L}\gg$ 1) at speeds above the gravitational escape speed from either planet (Hsu et al., 2012). Another is that the B ring of Saturn displays transient, 104 km scale, quasi-radial dust structures known as spokes, whose abundance is seasonally modulated by the ring illumination but also by Saturn’s rotating magnetic field (Smith et al. (1981), Porco & Danielson (1982)). The origin of spokes is still not fully understood, but it is clear that they consist of micron-sized dust particles briefly elevated above the Saturn ring plane and with motions that reflect the combined influence of local electrostatic and magnetic forces, as well as planetary gravity (Horányi et al., 2004).
@@ 8 Torques
Most asteroids and cometary nuclei are irregular in shape. On such objects, the non-gravitational forces described above generally do not pass through the center of mass, resulting in a torque. Left unchecked, the torque will drive inexorably towards rotational instability, where the centripetal forces exceed the gravitational and cohesive forces holding the body together. A strengthless oblate ellipsoid of density $\rho$ and with equatorial axes $a$ = $b$ and polar axis $c\leq a$ has critical period
$\tau_{crit}=\left(\frac{3\pi}{G\rho}\right)^{1/2}\left(\frac{a}{c}\right)^{1/2}.$ (33)
For a sphere ($a=c$) of density $\rho=10^{3}$ kg m-3, Equation 33 gives $\tau_{crit}$ = 3.3 hour. A highly oblate body of the same density but with $a/c$ = 2 would have a critical period 21/2 times longer, or 4.7 hours. A majority of the measured rotation periods of solar system bodies are within a factor of a few of $\tau_{crit}$ (Warner et al., 2009), suggesting the importance of centripetal forces. However, it must be admitted that observational biases against the measurement of periods either much shorter or much longer than a few hours are strong in most existing datasets.
@@ 8.1 Sublimation Torque
The magnitude of the sublimation torque on a comet is equal to the momentum lost per second in sublimated material multiplied by the moment arm, which we write as $k_{T}a$, where $a$ is the nucleus radius and 0 $\leq k_{T}\leq 1$ is a dimensionless multiplier. $k_{T}$ = 0 corresponds to perfectly isotropic sublimation with no net torque. $k_{T}$ = 1 corresponds to perfectly collimated ejection tangential to the surface of the body. Then, with a mass loss rate $\mu m_{H}Q_{g}$ [kg s-1] and an outflow speed $V_{th}$, the magnitude of the torque is $T=\mu m_{H}Q_{g}k_{T}V_{th}a$. The timescale for the torque to change the spin is then (Jewitt, 2021)
@@
The rotational lightcurves of some comets have enabled the measurement of spin changes (e.g., Kokotanekova et al. (2018)), allowing $\tau_{s}$ to be directly estimated. With additional observational constraints on $\rho$, $a$, $Q_{g}$ and $P$, and with the use of Equation 34, the dimensionless moment arm, $k_{T}$, can be estimated (Jewitt, 2021). The median value is $k_{T}$ = 0.007, meaning that only 0.7% of the outflow momentum is used to torque the nucleus. Even this tiny fraction is sufficient to quickly modify the spins of small comets.
@@
A necessary condition for Equation 34 to remain valid is the persistence of ice at or near the physical surface of the nucleus for timescales $\geq\tau_{s}$. Near surface ice persists because the speed with which the ice sublimation surface erodes into the nucleus exceeds the speed with which heat conducts into the interior, causing fresh ice to be continually excavated. The particular problem is that sublimation exposes refractory particles too large to be ejected by gas drag. These should eventually clog the surface, producing a “rubble mantle” and inhibiting or shutting down further sublimation. How this works in detail is not known, even after several years of in-situ investigation of the nucleus of 67P/Churyumov-Gerasimenko by the Rosetta spacecraft (Attree et al., 2023). Orbital evolution may play a role, particularly when the perihelion distance migrates to smaller values, leading to higher temperatures and sublimation fluxes. In the case of the active asteroids, ice is exposed only intermittently, perhaps in response to occasional impacts that clear a surface refractory mantle. The duty cycle (ratio of the time spent in sublimation to the total elapsed time) is less than $10^{-4}$ or even $10^{-5}$, allowing ice to persist for long times but rendering Equation 35 inapplicable to these objects.
@@ 8.2 YORP Torque
A major limitation to the order-of-magnitude estimates of the torque timescales, both for sublimation (Equation 35) and for YORP (Equation 37), concerns the stability of the torque. In comets, the magnitude and direction of the outgassing torque depend on the surface distribution and angular dependence of the outgassing. As the surface erodes, we expect the spatial and angular distribution of outgassing sources to change, altering the torque. For most comets, we possess few or no observational constraints on the areal distribution of sources or their evolution. Indeed, in-situ measurements from 67P/Churyumov-Gerasimenko show the difficulty in modeling this process even given the most detailed data (Attree et al., 2023). Likewise, the YORP torque is highly sensitive to the surface shape and texture and can change in magnitude and even direction in response to minor surface changes (Statler, 2009). Empirical but indirect evidence for this comes from asteroid (162173) Ryugu ($a\sim$ 0.5 km) which has the characteristic diamond shape indicative of rotational instability but a current rotation period near 7.6 hours. The rotation of Ryugu may have slowed in response to a changing YORP torque, leaving its equatorial ridge as evidence of its previously rapid spin. Spin evolution in the presence of a changing torque may be more akin to a random walk process than to the steady change implicit in Equations 35 and 37 (Bottke et al. (2015), Statler (2015)). The relevant spin-change timescales would then be much longer than estimated here. Unfortunately, we possess too little information to address this problem with any confidence.
@@ 8.3 Examples:
    Figure 17: Two active asteroids showing episodic ejections likely due to YORP-driven rotational instability. Each tail corresponds to a particular synchrone; the tail position angle is a measure of the date of ejection. 311P/Panstarrs (top) exhibited nine ejections over about nine months (six visible on 2013 Sep 10 plus three later) (Jewitt et al., 2013) while (6478) Gault displayed three tails in 2019 (Luu et al., 2021).
@@ 8.4 Summary
where $\sigma=5.67\times 10^{-8}$ W m-2 K-4 is the Stefan-Boltzmann radiation constant and we assume emissivity $\varepsilon$ = 1. The order-of-magnitude cooling timescale is $\tau_{c}\sim H/(dH/dt)$, so that we can use Equations 3 and 4 to write the thermal parameter, $\Theta=\tau_{c}/P$ (c.f., Spencer et al. (1989))
$\Theta=\frac{\rho c_{p}}{\sigma T^{3}}\left(\frac{\kappa}{P}\right)^{1/2}.$ (5)
Asteroids with very rapid rotation ($\Theta\gg$ 1) retain their heat through the night and should remain nearly longitudinally isothermal, producing little net diurnal Yarkovsky force. Similarly, asteroids with $\Theta\ll$ 1 will experience little diurnal Yarkovsky drift because rotation is too slow to carry the diurnal temperature maximum away from the midday meridian. Intuitively, on the other hand, asteroids with $\Theta\lesssim$ 1 should retain heat into the afternoon but lose it by the morning and so will experience a net force from asymmetric radiation [^5] With more effort than is warranted here, it can be shown that the effect is proportional to $\Theta/(1+2\Theta+2\Theta^{2})$, a function that is maximized for $\Theta\sim$ 0.7 (Rubincam, 1995)..
As a rough example, consider an asteroid with a very porous surface regolith having $\kappa$ = 10-9 m2 s-1 and with nominal $\rho=10^{3}$ kg m-3, $c_{p}=10^{3}$ J K-1 kg-1, $P=2\times 10^{4}$ s (i.e., about 6 hours) and orbiting near 1 au, where $T\sim$ 300 K. Substitution gives skin depth $\ell\sim$ 5 mm, $\Theta\sim$ 0.15, and the peak heat of noon will be retained for about 1/6th of a rotation, corresponding to a lag angle $\theta\sim$60°. The Yarkovsky force on a small asteroid with these parameters is likely to be significant. By comparison, an asteroid having the same properties but consisting of solid rock, for which $\kappa=10^{-6}$ m2 s-1 would have $\ell\sim$ 15 cm, $\Theta\sim$ 4.5 and, by virtue of taking several rotations in order to cool, would experience minimal diurnal temperature variation and reduced Yarkovsky acceleration.
@@
I thank Marco Fenucci for providing a digital table of his Yarkovsky drift measurements, Pedro Lacerda, Jane Luu, Joe Masiero, Darryl Seligman, David Vokrouhlicky, Emerson Whittaker and an anonymous referee for comments.